a2+2a-35 /12x devided by ax-3x/a2 + 4a -21


Found 2 solutions by RAY100, lazi:
Answer by RAY100(1637)

{(a^2 +2a -35) / 12x } / { ax-3x)/(a^2 +4a -21)
.
invert amd multiply last set of terms
.
{(a^2 +2a-35)/ 12x} * { (a^2 +4a -21) / (ax-3x)}
.
factor
.
{ (a+7)(a-5)/12x} * { ( a+7)(a-3)/x(a-3) }
.
simplify by canceling like terms
.
{(a+7)(a-5)(a+7)} / 12x^2 }
.
{ (a+7)^2 * (a-5) / 12x^2 }
.
This could be expanded, but it would not add to clarity
.
{a^3 +93a^2 -441a -245}/12x^2

Answer by lazi(15)

i'm assuming that this is what you meant to type
(a^2 + 2a - 35)/(12x) divided by (ax - 3x)/(a^2 + 4a - 21)

a^2 + 2a - 35 = (a + 7)(a - 5)
ax - 3x = x(a - 3)
a^2 + 4a - 21 = (a + 7)(a - 3)
when you have a fraction and you try to divide it to another fraction you actually multiplying its reciprical (ie 1/2 divide by 1/4 = 1/2 multiply by 4/1)
disregard the dots.
[(a+7)(a-5)]....[(a+7)(a-3)]....(a+7)^2 (a-5)
------------ x ------------ = -------------
(12x).....................x(a-3)................12x^2